OpenStudy (anonymous):

Solve square root of 2x plus 15 equals square root of 5

OpenStudy (zepp):

\(\sqrt{2x}+15=\sqrt{5}\) What would you get if you move 15 to the right?

OpenStudy (anonymous):

\( \sqrt{2x+15}=\sqrt{5}\) \(2x+15=5\) \(2x=-10\) \(x=-5\)

OpenStudy (zepp):

Umm, look like I misunderstood the question lol